A 216 sector of a circle of radius 5cm is bent to form a cone .what is the base radius. What is the vertical angle. What is the curved surface area
1
Expert's answer
2020-04-27T18:23:46-0400

Radius of the circle = "R=5cm," angle of the sector = "\theta=216\degree."

Length of the circular arc cut off from the circle 

"L=2\pi R\times{\theta \over 360\degree}=2\pi(5){216\degree \over 360\degree}=6\pi(cm)"

When the sector is cut and its bounding radii is bent to form a cone, slant height of the cone, "l=R=5cm."

Let "r"  and "h" be the radius and height of the cone formed.

Circumference of the base of the cone "=2\pi r=6\pi."


"r=3cm."

The base radius is "3cm."

Height of the cone


"h=\sqrt{l^2-r^2}=\sqrt{5^2-3^2}=4(cm)"

Let "\varphi" be a vertical angle. Then


"\sin(\varphi/2)={r \over l}={3\over 5}=0.6"

"\varphi=2\arcsin(0.6)\approx73.74\degree"

Find the curved surface area


"\pi rl=\pi(3)(5)=15\pi (cm^2)\approx47.12(cm^2)"